Inventory Costing Methods VanderMeer Inc. reported the following information for the month of October: Inventory, October 1 5I need the two blank boxes answered with the solutions
Answer #1

Units in ending inventory = Units available for sale - Units sold = (58+49+70+39) - 140 = 76 units

Cost of Goods Sold Ending inventory
Lifo

= (39*24)+(70*22)+(31*21)

= 3127

= (58*20)+(76-58)*21

= 1538

.

Comment if you face any issues